Parth (parthkohli):

@AravindG \(\sqrt{1} = \pm 1\) disprove.

OpenStudy (aravindg):

Ok first of all Theory part

OpenStudy (dumbsearch2):

Eh, guise. This is a continuation of http://openstudy.com/study#/updates/51a721e1e4b08f9e59e230bf in some respects. Just so you know, strangers xD

OpenStudy (dumbsearch2):

Let's kick back and watch the greenies argue

OpenStudy (aravindg):

We have 2 types of square roots ..Positive and negative .Unless stated explicitly we always take the positive square root.The square root function ALWAYS has a set of real positive numbers as range .Hence we denote the positive and negative square roots as : \(+\sqrt{x}\) and \(-\sqrt{x}\)

Parth (parthkohli):

Actually, we need to get back to the definition of a square root. If \(\sqrt{x} = n\), then the equation \(x = n^2 \) is satisfied. That is the definition. Does \(1 = (-1)^2\) satisfy? In that case, \(-1 = \sqrt{1}\). So is the case with \(n = 1\).

OpenStudy (aravindg):

The solutions for the equality \[x^2=1\] are 1 and -1 But notice how we get that : First root which is positive is \(+\sqrt{1}=1\) and the other negative root is \(-\sqrt{1}\)

OpenStudy (aravindg):

Notice that square root function always remains positive in the process

Parth (parthkohli):

Also, my question clearly on the top does not mention \(+\sqrt{1} = 1\) or \(-\sqrt{1} = 1\), but says \(\sqrt{1} = \pm1\) without any sign.

Parth (parthkohli):

Uh, actually, you are talking about \(-\sqrt{1}\) and \(+\sqrt{1}\). But what exactly \(\mathcal{is}\) \(\sqrt{1}\)?

OpenStudy (aravindg):

yes I talked abt them because they are the roots for x^2=1

Parth (parthkohli):

And what is your point that you are making?

OpenStudy (aravindg):

The square root function always remains positive

Parth (parthkohli):

Why does my use of the definition not convince you?

Parth (parthkohli):

You may be talking about the principal square root, but that has nothing to do with square root. They both are similar but different functions.

OpenStudy (aravindg):

Okay first I will provide you the standard definition \[\sqrt{x^2}=|x|\]

OpenStudy (aravindg):

This is how square root function is defined

OpenStudy (aravindg):

Notice the modulus sign on right ..That is why the range of it is positive real numbers

OpenStudy (aravindg):



OpenStudy (aravindg):

This is how it would look graphically

OpenStudy (aravindg):

I was drawing a rough diagram for square root function @dumbsearch2

Parth (parthkohli):

Are you sure that's the definition of a square root?

OpenStudy (aravindg):

yes 100%

OpenStudy (dumbsearch2):

Guys............ http://en.wikipedia.org/wiki/Square_root

OpenStudy (sasogeek):

okay to end all this shinanigans going on, let's just say that \(\huge \sqrt{x^2}=\pm x \) this is because, \(\huge +x^2=-x^2\) and both are always non-negative numbers. want proof? \(\huge \sqrt{4}=\pm2 \) \(\huge -2^2=4, \ and \ 2^2=4 \) everybody happy now?

Parth (parthkohli):

In mathematics, a square root of a number \(a\) is a number \(y\) such that \(y^2 = a\) :-P

Parth (parthkohli):

Let \(a = 1\) and \(y = -1\). Why is it that you think \(1 \ne (-1)^2\)?

OpenStudy (dumbsearch2):

"shinanigans" @sasogeek LOL! 1. shinanigans is spelled wrong. 2. Their debate is hardly shenanigans, it's just a real interesting discussion that fascinates our curious minds :D

OpenStudy (aravindg):

@ParthKohli Wolf agrees http://www.wolframalpha.com/input/?i=sqrt%7B1%7D

OpenStudy (sasogeek):

she or shi, u still got it :) it's a free world my friend xD

OpenStudy (dumbsearch2):

I was always taught that square rooting one made one. IDK though, cause I'm a n00b, just I learned it that way :)

Parth (parthkohli):

lol -- scroll down and look at "all 2nd roots of 1"

OpenStudy (dumbsearch2):

Wow. Parth was so right, the whole while. WOW! :D

OpenStudy (aravindg):

@dumbsearch2 no he isnt :/

OpenStudy (dumbsearch2):

@AravindG I know it's hard to admit :/

Parth (parthkohli):

Why are you dragging in the principal root all the time? What you said is just a property of square roots. It's not the definition!

OpenStudy (aravindg):

@saifoo.khan had the same misconception But we convinced him in a post before .

Parth (parthkohli):

@AravindG What was saif's question?

OpenStudy (aravindg):

No it didnt

OpenStudy (dumbsearch2):

Look, Wolfram|Alpha agreed with @ParthKohli. I've used their API before, they're extremely accurate.

OpenStudy (aravindg):

Maybe @.Sam. can assist me here

Parth (parthkohli):

What you're doing is actually neglecting a root. \(x^2 - x = 0 \Rightarrow x - 1 = 0 \Rightarrow x = 1\) is an example of neglecting a root.

Parth (parthkohli):

@.Sam.

Parth (parthkohli):

@experimentX

OpenStudy (dumbsearch2):

@dumbsearch2

OpenStudy (dumbsearch2):

:D

Parth (parthkohli):

What did Saifoo ask exactly?

OpenStudy (aravindg):

I told you we never neglected the roots The are 2 roots : \(+\sqrt{1}\) and \(-\sqrt{1}\)

Parth (parthkohli):

\(\sqrt{1}\) is like writing down \(\sin(\pi) \) for an answer. You actually want to write the numerical value.

OpenStudy (aravindg):

Gimme some time..let me think of how to explain it in a simpler way

Parth (parthkohli):

\[x^2 = 1 \Rightarrow x^2 - 1 = 0 \Rightarrow (x + 1)(x - 1) = 0 \Rightarrow x = \pm 1\]

OpenStudy (aravindg):

yes completely right ^

OpenStudy (aravindg):

and how we got it ? x=\(+\sqrt{1} \)and \(-\sqrt{1}\)

Parth (parthkohli):

WHY ARE YOU INCLUDING \(+\sqrt{1}\) AND \(-\sqrt1{}\)???????? I AM SIMPLY ASKING IF YOU CAN DISPROVE \(\sqrt1 = \pm 1 \)

OpenStudy (dumbsearch2):

Wow. Caps. Stay nice, buddies :P

Parth (parthkohli):

lol, you got the opposite. The domain is \(x \ge 0\) and the range is any real number.

OpenStudy (dumbsearch2):

Wow! @experimentX is dropping his 2 cents! :)